find dimensions of a triangular shape given lenght and a right triangle 12 inch legs
Answer #1

SOLUTION;

ABC is right angle triangle

Where A angle is right angle

And according to question c= 12 inches

C= hypotenuse

Suppose other two side is x,

So, a=b=x

Now,

c^{2}=a^{2}+b^{2}

(12)^{2}=x^{2}+x^{2}

(12)^{2}=2x^{2}

x= \sqrt{72} inches

Now, area of triangle is

A= \frac{1}{2}(base)(height)

A= \frac{1}{2}(x)(x)

A= \frac{1}{2}(x^{2})

A= \frac{1}{2}(72)

A=36 inch​​​2

Dimensions of triangle is

12 inch, \sqrt{72} inch, \sqrt{72} inch,